OpenStudy (anonymous):

frank has devised a formula for his catering buisness that calculates the number of meatballs he needs to prepare . the formula is m= 4a + 2c , where C equals the number of children , m = number of meatballs , and A = the number of adults .How many meatballs are required for a party of 25 adults and 5 children ?

OpenStudy (anonymous):

The problem only requires that you insert the variables where they are needed. m = 4(25) + 2(5) = 100 +10 = 110 Since the problem is a word problem , you may have to answer in a complete sentence. Answer: 110 meatballs are required for the party.